OpenStudy (anonymous):

Simplify the expression \[x^4+x^3-x^2+x-2\]

OpenStudy (anonymous):

I don't think you can

OpenStudy (anonymous):

The answer is: \[(x-1)(x+2)(x^2+1)\]

OpenStudy (anonymous):

so the question is to actually factor...

OpenStudy (anonymous):

just notice \[x^4+x^3-x^2+x-2=x^4-x^2+x^3-1+x-1\]

OpenStudy (anonymous):

do you know synthetic division?

OpenStudy (anonymous):

Yes I do... but that wouldn't buy me much time in a final exam...

OpenStudy (anonymous):

What do I do after that @mukushla

OpenStudy (anonymous):

Oh... I think I see, I'll try it again..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok, I lied... it's getting more complicated... I don't see it anymore...

OpenStudy (anonymous):

\[x^4-x^2=x^2(x-1)(x+1)\\ x^3-1=(x-1)(x^2+x+1)\] then x-1 is a factor of it

OpenStudy (anonymous):

I did that but I don't see how they come together to give me my final answer...

OpenStudy (anonymous):

now u have \[(x-1)(x^2(x+1)+x^2+x+2)=(x-1)(x^2(x+2)+x+2)\] then what will u get?

OpenStudy (anonymous):

No... sorry I still don't see it and where'd you get the 2 from.

OpenStudy (anonymous):

\[x^4−x^2+x^3−1+x−1=x^2(x-1)(x+1)+(x-1)(x^2+x+1)+x-1\\=(x-1)(x^2(x+1)+(x^2+x+1)+1)\]

OpenStudy (anonymous):

I still don't see how I can get my final answer...

OpenStudy (anonymous):

is that ok till there?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[=(x-1)(x^2(x+1)+x^2+x+2)=(x-1)(x^2(x+2)+x+2)=(x-1)(x+2)(x^2+1)\]

OpenStudy (anonymous):

I see how you got the second line but I don't see how you got the line after that...

OpenStudy (anonymous):

(x+2) is one of the factors of mid line x^2(x+2)+x+2=(x+2)(x^2+1)

OpenStudy (anonymous):

Oh... wow... why am I forgetting the simple stuff.. Okay, thanks so much @mukushla

OpenStudy (anonymous):

your welcome sorry my english is not well and i cant explain well

OpenStudy (anonymous):

That's okay! You explained it terrifically!